# A large Deviation Principle for the Normalized Excursion of $\alpha$-Stable Lévy Processes without Negative Jump

### v2

@@ 6.   Consequences of the LDP for stable excursions
In this section, we prove the remaining statements: 1.4, Corollaries 1.5 and 1.6, and Proposition 1.3.
6.1.  Proof of 1.4
We follow the approach of Fill and Janson [11] closely. First, a direct consequence of the fact that $I_{\mathbbm{e}}$ is a good rate function is the following.
@@
The argument for the proof of 1.4 is the same as [11, p.415]. We apply the contraction principle ([15, Theorem 27.11], [9, Theorem 4.2.1]) to the continuous functional $\Phi:D_{\mathrm{ex}}[0,1]\rightarrow\mathbb{R}_{+}$. This entails that $\varepsilon X=\Phi(\varepsilon\mathbbm{e})$ satisfies an LDP in $[0,\infty)$ with good rate function whose value at $x>0$ is given by
@@
Taking $A=(1,\infty)$ and $\varepsilon=x^{-1}$ in the definition of an LDP proves (1.8). Finally, (1.9) and (1.10) follow from (1.8) by [14, Theorem 4.5]. This concludes the proof of 1.4. ∎
@@ 6.2.  Applications
Corollaries 1.5 and 1.6 are obtained by applying 1.4 to the area and supremum functionals, which are both positive-homogeneous and continuous for the M1 topology.
@@ 6.3.  Negative results for the Skorokhod J1 topology
Here we give a sketch of proof for 1.3. The idea is that it is costless for the process $\varepsilon\mathbbm{e}$ to make macroscopic jumps within small time intervals, which prevents it from being concentrated in J1-compact sets. Fix $\delta>0$. We note that
$\mathbb{P}\left(\varepsilon\mathbbm{e}_{\delta}\in[1,2],\varepsilon\mathbbm{e}_{2\delta}\in[3,4]\right)=\alpha\Gamma\left(1-\frac{1}{\alpha}\right)\int_{1/\varepsilon}^{2/\varepsilon}\,\mathrm{d}x_{1}\,j_{\delta}(x_{1})\int_{3/\varepsilon}^{4/\varepsilon}\,\mathrm{d}x_{2}\,p_{\delta}^{(0,\infty)}(x_{1},x_{2})q_{x_{2}}(1-2\delta)\\ \geq C(\delta)\varepsilon^{2\alpha+2}\exp(-c_{\alpha}(3/\varepsilon(1-2\delta))^{\alpha^{\prime}})\,,$ (6.1)
@@ 7.   LDP for the $\alpha$-stable Lévy bridge
In this section, we adapt the proof of the LDP for the normalized excursion in order to get an LDP for the Lévy bridge. Roughly speaking the process $\mathbbm{b}^{(a)}$, called the $(0,0)\rightarrow(1,a)$ bridge, is obtained by conditioning $L$ to be equal to $a$ at time $1$. This is obviously a degenerate conditioning; however, it can be obtained by performing a space-time $h$-transform with respect to the function $\frac{p_{1-t}(a-L_{t})}{p_{1}(a)}$ (see, for instance, [19]). This means that the law of $\mathbbm{b}^{(a)}$ may be defined by
$\mathbb{P}^{\textrm{br}}(A)\coloneqq\mathbb{E}\left[\frac{p_{1-t}(a-L_{t})}{p_{1}(a)}\mathbbm{1}_{A}\right],\quad\forall A\in\mathcal{F}_{t}\,,\quad t\in[0,1).$ (7.1)
See [8] or [3, Chapter VIII] for a rigorous construction.
@@ 7.1.  Large deviations for the finite-dimensional marginal distributions
The proof of this proposition is similar to that of Proposition 3.1, but is technically simpler and we only explain where the argument differs. It is not difficult to check that $J_{\mathbbm{b},a}$ is a good rate function, so [29, Lemma 5] applies. We use this in the same way as in the proof of Proposition 3.1, using the expression (2.6) for the marginal laws of the bridge, and the bounds (3.3) for the stable transition densities. The term $-(a_{-})^{\alpha^{\prime}}$ in the definition of $J_{\mathbbm{b},a}$ arises from the contribution of the density $p_{1}(a_{\varepsilon})$ in the denominator of (2.6): by (3.3) and (3.4), we have
@@ 7.2.  Exponential tightness for the stable bridge
7.2 is a direct consequence of the tightness criterion in $(\mathbb{D}[0,1],\mathrm{dist})$ from 2.2 and the following lemma.
@@
Splitting the expectation into five terms as at the beginning of the proof of 4.1, we get the following bound
$\mathbb{E}\left[\exp\big{(}\lambda M(\mathbbm{b}^{(a_{\varepsilon})}_{s},\mathbbm{b}^{(a_{\varepsilon})}_{t},\mathbbm{b}^{(a_{\varepsilon})}_{u})\big{)}\right]\leq 2\mathbb{E}\left[e^{\lambda(\mathbbm{b}^{(a_{\varepsilon})}_{s}-\mathbbm{b}^{(a_{\varepsilon})}_{t})}\mathbbm{1}_{\{\mathbbm{b}^{(a_{\varepsilon})}_{t}\leq\mathbbm{b}^{(a_{\varepsilon})}_{s}\}}\right]\\ +2\mathbb{E}\left[e^{\lambda(\mathbbm{b}^{(a_{\varepsilon})}_{t}-\mathbbm{b}^{(a_{\varepsilon})}_{u})}\mathbbm{1}_{\{\mathbbm{b}^{(a_{\varepsilon})}_{u}\leq\mathbbm{b}^{(a_{\varepsilon})}_{t}\}}\right]+1.$
@@
7.3.  Proof of 1.7
We may finally prove 1.7 using 7.1 and 7.2. The scheme of proof is exactly the same as that in Section 5, and combines the exponential tightness in $(\mathbb{D}[0,1],\mathrm{dist})$ with an LDP in the weak topology $(\mathbb{D}[0,1],\mathcal{W})$, similar to Proposition 5.1. Therefore, we will give a brief account, only pointing out the places where the formulas differ.
We can easily adapt the proof of 5.2 to get the following proposition. For a subdivision $\sigma=(t_{1},\ldots,t_{n})$ and $f\in\mathbb{D}[0,1]$, define $I_{\mathbbm{b},a}^{\sigma}(f)=J_{\mathbbm{b},a}^{\sigma}(f(t_{1}),\ldots,f(t_{n}))$.
@@
$M_{\mathbbm{b},a}(f)\coloneqq\sup_{\sigma\in\mathfrak{S}}I_{\mathbbm{b},a}^{\sigma}(f)<\infty.$
@@
Then, analogs of Lemmas 5.3, 5.4, 5.6 and 5.6 hold true with $I_{\mathbbm{b},a}$ in place of $I_{\mathbbm{e}}$, with exactly the same proofs. This ends the proof of 1.7.

### v3

@@ 6. Consequences of the LDP for stable excursions
In this section, we prove the remaining statements: Theorem 1.4, Corollaries 1.5 and 1.6, and Proposition 1.3.
6.1. Proof of Theorem 1.4
We follow the approach of Fill and Janson (2009) closely. First, a direct consequence of the fact that $I_{\mathbbm{e}}$ is a good rate function is the following.
@@
The argument for the proof of Theorem 1.4 is the same as (Fill and Janson, 2009, p.415). We apply the contraction principle ((Kallenberg, 2002, Theorem 27.11), (Dembo and Zeitouni, 1998, Theorem 4.2.1)) to the continuous functional $\Phi:D_{\mathrm{ex}}[0,1]\rightarrow\mathbb{R}_{+}$. This entails that $\varepsilon X=\Phi(\varepsilon\mathbbm{e})$ satisfies an LDP in $[0,\infty)$ with good rate function whose value at $x>0$ is given by
@@
Taking $A=(1,\infty)$ and $\varepsilon=x^{-1}$ in the definition of an LDP proves (1.8). Finally, (1.9) and (1.10) follow from (1.8) by (Janson and Chassaing, 2004, Theorem 4.5). This concludes the proof of Theorem 1.4. ∎
@@ 6.2. Applications
Corollaries 1.5 and 1.6 are obtained by applying Theorem 1.4 to the area and supremum functionals, which are both positive-homogeneous and continuous for the M1 topology.
@@ 6.3. Negative results for the Skorokhod J1 topology
Here we give a sketch of proof for Proposition 1.3. The idea is that it is costless for the process $\varepsilon\mathbbm{e}$ to make macroscopic jumps within small time intervals, which prevents it from being concentrated in J1-compact sets. Fix $\delta>0$. We note that
$\mathbb{P}\left(\varepsilon\mathbbm{e}_{\delta}\in[1,2],\varepsilon\mathbbm{e}_{2\delta}\in[3,4]\right)$ $=\alpha\Gamma\left(1-\frac{1}{\alpha}\right)\int_{1/\varepsilon}^{2/\varepsilon}\,\mathrm{d}x_{1}\,j_{\delta}(x_{1})\int_{3/\varepsilon}^{4/\varepsilon}\,\mathrm{d}x_{2}\,p_{\delta}^{(0,\infty)}(x_{1},x_{2})q_{x_{2}}(1-2\delta)$
$\geq C(\delta)\varepsilon^{2\alpha+2}\exp(-c_{\alpha}(3/\varepsilon(1-2\delta))^{\alpha^{\prime}})\,,$ (6.1)
@@ 7. LDP for the $\alpha$-stable Lévy bridge
In this section, we adapt the proof of the LDP for the normalized excursion in order to get an LDP for the Lévy bridge. Roughly speaking the process $\mathbbm{b}^{(a)}$, called the $(0,0)\rightarrow(1,a)$ bridge, is obtained by conditioning $L$ to be equal to $a$ at time $1$. This is obviously a degenerate conditioning; however, it can be obtained by performing a space-time $h$-transform with respect to the function $\frac{p_{1-t}(a-L_{t})}{p_{1}(a)}$ (see, for instance, Liggett (1968)). This means that the law of $\mathbbm{b}^{(a)}$ may be defined by
$\mathbb{P}^{\textrm{br}}(A)=\mathbb{E}\left[\frac{p_{1-t}(a-L_{t})}{p_{1}(a)}\mathbbm{1}_{A}\right],\quad\forall A\in\mathcal{F}_{t}\,,\quad t\in[0,1).$ (7.1)
See Chaumont (1997) or (Bertoin, 1996, Chapter VIII) for a rigorous construction.
@@ 7.1. Large deviations for the finite-dimensional marginal distributions
The proof of this proposition is similar to that of Proposition 3.1, but is technically simpler and we only explain where the argument differs. It is not difficult to check that $J_{\mathbbm{b},a}$ is a good rate function, so (Serlet, 1997, Lemma 5) applies. We use this in the same way as in the proof of Proposition 3.1, using the expression (2.6) for the marginal laws of the bridge, and the bounds (3.3) for the stable transition densities. The term $-(a_{-})^{\alpha^{\prime}}$ in the definition of $J_{\mathbbm{b},a}$ arises from the contribution of the density $p_{1}(a_{\varepsilon})$ in the denominator of (2.6): by (3.3) and (3.4), we have
@@ 7.2. Exponential tightness for the stable bridge
Proposition 7.2 is a direct consequence of the tightness criterion in $(\mathbb{D}[0,1],\mathrm{dist})$ from Theorem 2.2 and the following lemma.
@@
Splitting the expectation into five terms as at the beginning of the proof of Proposition 4.1, we get the following bound
$\mathbb{E}\left[\exp\big{(}\lambda M(\mathbbm{b}^{(a_{\varepsilon})}_{s},\mathbbm{b}^{(a_{\varepsilon})}_{t},\mathbbm{b}^{(a_{\varepsilon})}_{u})\big{)}\right]\\ \leq 2\mathbb{E}\left[e^{\lambda(\mathbbm{b}^{(a_{\varepsilon})}_{s}-\mathbbm{b}^{(a_{\varepsilon})}_{t})}\mathbbm{1}_{\{\mathbbm{b}^{(a_{\varepsilon})}_{t}\leq\mathbbm{b}^{(a_{\varepsilon})}_{s}\}}\right]+2\mathbb{E}\left[e^{\lambda(\mathbbm{b}^{(a_{\varepsilon})}_{t}-\mathbbm{b}^{(a_{\varepsilon})}_{u})}\mathbbm{1}_{\{\mathbbm{b}^{(a_{\varepsilon})}_{u}\leq\mathbbm{b}^{(a_{\varepsilon})}_{t}\}}\right]+1.$
@@
7.3. Proof of Theorem 1.7
We may finally prove Theorem 1.7 using Propositions 7.1 and 7.2. The scheme of proof is exactly the same as that in Section 5, and combines the exponential tightness in $(\mathbb{D}[0,1],\mathrm{dist})$ with an LDP in the weak topology $(\mathbb{D}[0,1],\mathcal{W})$, similar to Proposition 5.1. Therefore, we will give a brief account, only pointing out the places where the formulas differ.
We can easily adapt the proof of Proposition 5.2 to get the following proposition. For a subdivision $\sigma=(t_{1},\ldots,t_{n})$ and $f\in\mathbb{D}[0,1]$, define $I_{\mathbbm{b},a}^{\sigma}(f)=J_{\mathbbm{b},a}^{\sigma}(f(t_{1}),\ldots,f(t_{n}))$.
@@
$M_{\mathbbm{b},a}(f)=\sup_{\sigma\in\mathfrak{S}}I_{\mathbbm{b},a}^{\sigma}(f)<\infty.$
@@
Then, analogs of Lemmas 5.3, 5.4, 5.5 and 5.6 hold true with $I_{\mathbbm{b},a}$ in place of $I_{\mathbbm{e}}$, with exactly the same proofs. This ends the proof of Theorem 1.7.
@@
Acknowledgements
Thanks are due to Loïc Chaumont for an interesting conversation around stable excursions. We also thank the anonymous referee for their careful reading of the paper, and for bringing a number of relevant references to our attention.